2 long parallel wires (in the plane of the page) each 1.0 m long, separated by 2.0 cm, each carry currents of 1.0 A from bottom towards the top of the page. (a) Calculate the magnetic field at the point P midway between the wires in the plane of the page. And (b) Calculate the magnetic field at the point Q 1.0 cm from the right of the wire 2. The point Q is in the plane of the page.
Answer #1

Solution:

THe magnetic field due to current carrying long wire B = \mu I/2\pi r

In the midway between two wires, the direction of magnetic field will add each other and will be in negative y direction.

So, Beff = \mu X1/2\pi (1/0.01 + 1/0.01)

= 2X10^-7X200 = 4X10^-5 Ans

In the second part, both the magnetic fields will oppose each other in direction. THe direction due to right wire will will be in positive y direction while the direction due to left wire will be negative y direction.

So, in this case Effective Magnetic Field Beff   = \mu X1/2\pi (1/0.01- 1/0.03)

= 2X10^-7X66 = 1.33X10^-5 Ans